OpenStudy (anonymous):

simplify the trigonometric expression. sin^2 theta/1-cos theta a. 1+cos theta b. son theta c. 1-sin theta/ cos theta d. 1+sin theta/cos theta

OpenStudy (imstuck):

Here you have to use a Pythagorean Identity and then do some factoring. If \[\sin ^{2}\theta + \cos ^{2}\theta= 1\]what does \[\sin ^{2}\theta \]equal? Solve that equation for sin^2 theta.

OpenStudy (imstuck):

Alrighty then!

OpenStudy (anonymous):

okay hang on let me look at this carefully

OpenStudy (anonymous):

i don't know what all those symbols mean that your using...